The inverse demand function is given by P = 2 – 0.13 Q.

What is the price elasticity of demand?

What is the price elasticity of demand when P = 1.2


1
Expert's answer
2021-03-04T15:05:53-0500

The price elasticy of demand is


"\epsilon=\dfrac{P}{Q}\cdot\dfrac{dQ}{dP}"

"P=2-0.13Q"

"Q=\dfrac{2-P}{0.13}"

"\dfrac{dQ}{dP}=-\dfrac{1}{0.13}"

"\epsilon=\dfrac{0.13P}{2-P}\cdot(-\dfrac{1}{0.13})"

"\epsilon=\dfrac{P}{P-2}"

"P=1.2"


"\epsilon=\dfrac{1.2}{1.2-2}=-1.5"

"\epsilon=-1.5"